OpenStudy (anonymous):

if (3,y) is on the line 3x-4y =10, what is y?

OpenStudy (anonymous):

3(3)-4y=10, solve for y and you got it

OpenStudy (anonymous):

replace \(x\) by \(3\) and solve for \(y\) that is, solve \[3\times 3-4y=10\] for \(y\)

OpenStudy (anonymous):

Usain and Blake.